Suppose a firm is made up of 58% equity, 7% preferred stock, and the remainder is debt. IF the cost to raise debt for this firm is 4%, the cost to raise preferred stock is 5%, and the cost to raise equity is 8%, what is the weighted average flotation cost?

A- 6.04%

B- 5.83%

C- 6.77%

D- 6.39%

E-7.16%

Please show the calculation steps so I can understand how to solve.
Thank you!
Answer #1

Weight of debt=100-(58+7)=35%

weighted average flotation cost=Respective costs*Respective weight

=(0.58*8)+(0.07*5)+(0.35*4)

which is equal to

=6.39%